# SIRD: A Sender-Informed, Receiver-Driven Datacenter Transport Protocol

### v4

@@ 4.3. Informed Overcommitment
1:Variables: $\bm{c_{r}}$: available credit for outbound messages to receiver r
2:procedure onCreditPacket($r,credit$)
3:    $c_{r}\leftarrow c_{r}+credit$
4:end procedure
5:procedure poll()
6:    $rcvrList\leftarrow activeReceivers.filter(c_{r}>0)$;
7:    $(r,dataPkt)\leftarrow{policySelect}(rcvrList)$
8:    $dataPkt.sird.csn\leftarrow(\sum{{}_{i}}(c_{i})\geq SThr)$
9:    $c_{r}\leftarrow c_{r}-dataPkt.size$
10:    ${send}(r,dataPkt)$
11:end procedure
4.4. Congestion Control Algorithm
Algorithm 1 describes the actions of a SIRD receiver. Assuming credit is available in the global bucket, the receiver tries to allocate it to an active sender (ln. 9). It first, selects one of the senders with available credit in the per-sender bucket (ln.10) based on policy (ln.11). The receiver sends a CREDIT packet (ln. 13) and reduces the available credit both in the global bucket, the per-sender bucket, and the remaining required credit for that message (ln. 14). Whenever a DATA packet arrives (ln. 2), if it is scheduled, the receiver replenishes credit in the global bucket (ln. 4) and the per-sender bucket (ln. 5). Then it executes the two independent AIMD control loops to adapt to a congested sender (ln. 6) and a congested core network (ln. 7), respectively. The receiver sets the size of the per-sender bucket as the minimum of the two values determined by the control loops (ln. 10).
@@ 4.4. Congestion Control Algorithm
Algorithm 2 describes the implementation of the sender-side algorithm for scheduled DATA packets. A host can send data to a receiver as long as it has available credit for said receiver. Congested senders mark the congested sender notification (sird.csn) bit if the total amount of accumulated credit exceeds $SThr$ (ln. 8). The sender can send unscheduled DATA packets at any point in time. SIRD senders naturally handle scenarios where a meaningful portion of uplink bandwidth is consumed by unscheduled packets because they limit credit accumulation and thus adapt the transmission rate of scheduled packets to the leftover bandwidth.
4.5. Packet Pacing and Scheduling
@@
SIRD can be configured to schedule for fairness e.g., by crediting messages in a round-robin manner, or for latency minimization, e.g., by crediting smaller messages first (SRPT), or to accommodate different tenant classes. SIRD implements policies at the receiver (ln .11), which is the primary enforcer, and at the sender (ln. 7). By minimizing queuing in the fabric, SIRD does not need to enforce policies there, simplifying the design. Regardless of which policy is configured at senders, SIRD allocates part of the uplink bandwidth fairly across active receivers, as to ensure a regular flow of congestion information between sender-receiver pairs.
@@
SIRD receivers pace credit transmission to match the downlink’s capacity for data packets. Pacing helps further reduce downlink queuing but is not needed for correctness.
4.6. Ethernet Priorities and Packet Loss
Whenever possible in the datacenter, CREDITREQ and CREDIT packets are sent at a higher network priority to further reduce RTT jitter. The unscheduled prefixes of messages $<UnschT$ also use it. Using one higher priority level for control and prefix packets has minimal impact on the overall traffic and should not affect other operations. SIRD does not require priorities to deliver high performance (§5.5).
Packet loss in SIRD will be very rare by design, but the protocol must nevertheless operate correctly in its presence. Ethernet packets may be dropped for reasons other than congestion, e.g., CRC errors or packet drops due to host, switch, or link-level restarts. SIRD employs Homa’s (Montazeri et al., 2018) retransmission design which uses timeouts at the receiver and the sender to detect packet loss.
5. Protocol Evaluation
@@
We use network simulation to evaluate SIRD in complex scenarios. We seek to answer questions in four areas: (1) How well does SIRD navigate the throughput-buffering-latency trade-off (§5.2)? (2) Is SIRD’s congestion response robust against high load pressure and adverse traffic patterns? Does it effectively deal with core congestion (§5.3)? (3) Can SIRD deliver messages with low and predictable latency (§5.4)? (4) How important is SIRD’s sender-informed design in maximizing utilization while minimizing buffering? How sensitive is SIRD to its configuration parameters (§5.5)?
5.1. Methodology
@@
Protocols: We compare SIRD to 5 baselines: DCTCP (Alizadeh et al., 2010), a widely deployed sender-driven scheme (Ghabashneh et al., 2022), Swift (Kumar et al., 2020), a state-of-the-art production sender-driven scheme, Homa (Montazeri et al., 2018), because of its near-optimal latency and its use of overcommitment, ExpressPass (Cho et al., 2017), as it employs a hop-by-hop approach to managing credit, and dcPIM (Cai et al., 2022), as a recent and unique point in the design space as it explicitly matches senders and receivers. We do not extend Homa with Aeolus (Hu et al., 2020) for the reasons discussed in (John K. Ousterhout, 2022a). Note that the Homa and dcPIM papers already include favorable comparisons to NDP (Handley et al., 2017), Aeolus (Hu et al., 2020), PIAS (Bai et al., 2015), pHost (Gao et al., 2015), and HPCC (Li et al., 2019), thus we did not include them in our evaluation.
@@
We implement SIRD on ns-2 (Altman and Jimenez, 2012), reuse the original ns-2 DCTCP implementation (using the same parameters as (Alizadeh et al., 2013), scaled to 100Gbps) and the original ns-2 ExpressPass implementation (Cho et al., 2018), and port the published Homa simulator (Montazeri, 2019) to ns-2 using the same parameters as in (Montazeri et al., 2018) scaled to 100Gbps. The published Homa simulator does not implement the incast optimization (Montazeri et al., 2018), which further relies on two-way messages. We use a Swift simulator, kindly made public (Abdous et al., 2021), and configure its delay parameters to achieve similar throughput to DCTCP, respecting the guidelines (Kumar et al., 2020). Finally, we use the published dcPIM simulator (Cai, 2023).
Table 2. Default simulation parameters for each protocol.
@@
Homa Same as (Montazeri et al., 2018) at 100Gbps (incl. priority split)
dcPIM Same as (Cai et al., 2022)
Table 2 lists protocol parameter values. DCTCP and Swift use pools of pre-established connections (40 for each host pair). SIRD approximates SRPT scheduling like Homa and dcPIM. SIRD, Homa, and dcPIM use packet spraying, DCTCP and Swift use ECMP. Homa uses 8 network priority levels, dcPIM 3, and SIRD 2 (§4.6). SIRD does not require priorities to operate well; we explore this in §5.5. The initial window of DCTCP and Swift is configured at $BDP$ like in (Alizadeh et al., 2013).
Topology: We simulate the two-tier leaf-spine topology used in previous work (Montazeri et al., 2018; Alizadeh et al., 2010, 2013; Gao et al., 2015; Cai et al., 2022) with 144 hosts, connected to 9 top-of-rack switches (16 hosts each) via 4 spine switches, with link speeds of 100Gbps to hosts and 400Gbps to spines; i.e., each switch has 3.2Tbps of bisection bandwidth. We simulate switches with infinite buffers, i.e., without packet drops (1) to avoid making methodologically complex assumptions as drop rates and thus latency and throughput are very sensitive to switch buffer sizes, organization, and configuration (Ghabashneh et al., 2022; John K. Ousterhout, 2022a) (see  §2.2); and (2) to study intended mode of operation of the protocols that leverage buffering to achieve high link utilization and operate best without drops. SIRD never uses more than a small fraction of the theoretical capacity and does not benefit from this setup.
Workload: As in prior work, each host operates as a client and a server, with one-way messages generated according to an open-loop Poisson distribution to uniformly random receivers (all-to-all). We simulate 3 workloads: (1) WKa: an aggregate of RPC sizes at a Google datacenter (Felderman, 2018); (2) WKb: a Hadoop workload at Facebook (Roy et al., 2015); (3) WKc: a web search application (Alizadeh et al., 2013). We select them to test over a wide range of mean message sizes of 3KB, 125KB, and 2.5MB respectively.
We simulate these 3 workloads on 3 traffic configurations, for a total of 9 points of comparison. The configurations are: (1) Balanced: The default configuration described above. We vary the applied load, which does not include protocol-dependent header overheads, from 25% to 95% of link capacity. (2) Core: Same as (1) but ToR-Spine links are 200Gbps (2-to-1 oversubscription). Due to uniform message target selection, $128/144\approx 89\%$ of messages travel via spines, turning the core into the bottleneck. We consequently reduce the load applied by hosts by ×$\times 0.89*2$ to reflect the network’s reduced capacity. This is not meant to reflect a permanent load distribution, but we hypothesize that it is possible transiently. (3) Incast: We use the methodology of (Li et al., 2019; Cai et al., 2022) and combine background traffic with overlay incast traffic: 30 random senders periodically send a 500KB message to a random receiver. Incast traffic represents 7% of the total load.
We report the goodput (rate of received application payload), total buffering in switches, and message slowdown, defined as the ratio between the measured and the minimum possible latency for each message. In the incast configuration, we exclude incast messages from slowdown results.
@@
Figure 4. Normalized goodput, queuing, and slowdown across all 9 configurations. Each metric is normalized based on the best-performing protocol for each metric and scenario. Normalized slowdown and buffering are capped at 10×10$\times$10 × and 200×200$\times$200 ×, respectively and higher values are plotted in the overflow area. X-axis jitter is added for visibility. See Table 4 for the data.
5.2. Performance Overview
@@
Figure 4 shows how protocols navigate the trade-offs between throughput, buffering, and latency by plotting their relative performance across all 9 workload-configuration combinations. The best-performing protocol on each of the 9 scenarios gets a score of $1.0$ (per metric) and the others are normalized to it, so that goodput is always $\leq 1.0$ whereas queuing and slowdown are always $\geq 1.0$. We report the highest achieved goodput and the highest peak queuing over all the levels of applied load. We report slowdown at 50% applied load which is a level most protocols can deliver in all scenarios. The following figures and calculations do not include cases where a protocol cannot satisfy a specific load level, or cannot stop network buffers from growing infinitely.
@@
Overall, SIRD is the only protocol that consistently achieves near-ideal scores across all metrics. Specifically, SIRD causes 12×12$\times$12 × less peak network queuing than Homa and achieves similar latency and goodput performance. SIRD outperforms dcPIM in message slowdown, and peak goodput and queuing by $46\%$, $9\%$, and $43\%$ respectively. ExpressPass causes practically zero queuing thanks to its hop-by-hop design, and 88% less than SIRD, but SIRD delivers 10×10$\times$10 × lower message slowdown and $26\%$ more goodput. SIRD never causes more than 0.8MB and 2.3MB in ToR queuing on receiver- and core-bottleneck scenarios respectively, which maps to 8% and 23% of the packet buffer size assuming 3.13 MB/Tbps (§2.2).
5.3. Congestion Response

### v5

@@ 4.3 Congestion Control Algorithm
Variables: $\bm{c_{r}}$: available credit for outbound messages to receiver r
1:procedure onCreditPacket($r,credit$)
2:    $c_{r}\leftarrow c_{r}+credit$
3:end procedure
4:procedure sendData()
5:    $rcvrList\leftarrow activeReceivers.filter(c_{r}>0)$;
6:    $(r,dataPkt)\leftarrow{policySelect}(rcvrList)$
7:    $dataPkt.sird.csn\leftarrow(\sum{{}_{i}}(c_{i})\geq SThr)$
8:    $c_{r}\leftarrow c_{r}-dataPkt.size$
9:    ${send}(r,dataPkt)$
10:end procedure
@@
Algorithm 2 describes the implementation of the sender-side algorithm for scheduled DATA packets. A host can send data to a receiver only if it has credit from said receiver. Congested senders mark the congested sender notification (sird.csn) bit if the total amount of accumulated credit exceeds $SThr$ (ln. 7). The sender can send unscheduled DATA packets at any point in time. It is worth noting that SIRD senders naturally handle scenarios where a meaningful portion of uplink bandwidth is consumed by unscheduled packets. Since senders limit credit accumulation by informing receivers to allocate less credit, the transmission rate of scheduled packets converges to the leftover bandwidth.
4.4 Other Design Concerns
@@
SIRD can be configured to schedule for fairness e.g., by crediting messages in a round-robin manner, or for latency minimization, e.g., by crediting smaller messages first (SRPT), or to accommodate different tenant classes. SIRD implements policies at the receiver (ln .10), which is the primary enforcer, and at the sender (ln. 6). By minimizing queuing in the fabric, SIRD does not need to enforce policies there, simplifying the design. Regardless of which policy is configured at senders, SIRD allocates part of the uplink bandwidth fairly across active receivers, as to ensure a regular flow of congestion information between sender-receiver pairs.
@@
SIRD receivers pace credit transmission to match their downlink’s capacity. Pacing improves message latency by reducing downlink queuing from scheduled packets even below the tight $B-BDP$ bound but is not needed for correctness.
If switch priority queues are available, CREDIT packets are sent over a higher priority lane to further reduce RTT jitter. The unscheduled prefixes of messages smaller than $UnschT$ also use this lane to bypass transient network queueing. SIRD does not require in-network priority queues to deliver high performance but sees tail latency benefits in some cases (§6.2.4).
Packet loss in SIRD will be very rare by design, but the protocol must still operate correctly in the presence of CRC errors or packet drops due to faults or restarts. SIRD employs Homa’s [56] retransmission design in which receivers infer loss when no new packets are received for an incomplete message after a period of a few milliseconds. Upon detecting the loss of a scheduled segment, receivers also reclaim the credit allocated for said segment. SIRD’s modest credit overcommitment (e.g., 50%) ensures that receivers can continue fully utilizing their bandwidth even in the improbable event where a sizeable chunk of packets is lost. Finally, SIRD incurs no performance penalties from out-of-order packet deliveries as it does not rely on packet order for loss detection and only delivers completed messages to applications.
5 Caladan Implementation
SIRD, with its message-level API and simple credit management design, can be implemented in NIC hardware or in software. The former eliminates PCIe latencies from the RTT, reducing the $BDP$, and allows for nanosecond-scale pacing and delivery of packets on the wire. We demonstrate that the latter is also possible, at 100Gbps, by implementing SIRD in 4300 SLOC as a UDP-based transport protocol in Caladan [29], a state-of-the-art kernel-bypass dataplane operating system and CPU scheduler. We chose Caladan because of its mature network stack and its lightweight green thread abstraction, suitable for building latency-sensitive applications. SIRD is open source and available on GitHub [1].
The SIRD stack follows an asymmetric design where dedicated spinning threads perform specialized tasks. On the receiver-side, each SIRD node has one dedicated thread that runs the control loops and implements downlink scheduling policies, like round-robin or SRPT, by transmitting paced CREDITs according to Algorithm 1. SIRD reduces network queuing by pacing credit at slightly less than the line rate, as proposed in Hull [9]. Dedicated receiver cores read from RX rings and adjust credits, ensuring the system can quickly replenish new credits for redistribution. Each dedicated receiver core has two RX rings, one for packets of scheduled messages and one for packets of unscheduled messages; this helps avoid drops of scheduled packets during high rates of unscheduled traffic. For efficiency, and to sustain 100Gbps line rates, memory allocation and data copying operations are removed from the credit replenishment path to a small set of dedicated threads which copy packets, reassemble messages from packets, which can be received out of order, and deliver messages to applications via inline callbacks.
The sender-side implementation also has a central thread, which accepts CREDIT packets and schedules the transmission of all DATA packets, whether these are unscheduled or scheduled. This thread implements Algorithm 2 which marks the $sird.csn$ bit of DATA packets by comparing the amount of accumulated credit to $SThr$. Like the central receiver thread, this thread paces itself to ensure that packets are not given to the NIC faster than it can consume them. The thread accepts new message requests via asynchronous shared-memory queues, where the application API returns error codes if resource limits are exceeded.
6 Evaluation
@@
We evaluate SIRD using our Caladan implementation and through simulation to answer the following questions: (1) Can a software implementation of SIRD deliver 100Gbps along with minimal queuing under incast (§6.1.1) and efficient credit management under outcast (§6.1.2)? (2) How well does SIRD navigate the throughput-buffering-latency trade-off in larger scale workloads compared to existing work (§6.2.1)? (3) Is SIRD’s congestion response robust at high load pressure and can it handle core congestion (§6.2.2)? (4) Can SIRD deliver messages with low latency in large scale workloads (§6.2.3)? (5) How important is SIRD’s sender-informed design in maximizing utilization with minimal buffering? How sensitive is SIRD to its parameters (§6.2.4)?
6.1 System Evaluation
We evaluate SIRD at a small scale using Cloudlab’s [27] sm110p machines equipped with 100Gbps ConnectX-6 DX NICs and Xeon 4314 CPUs, all located in the same rack. Our implementation uses 9KB jumbo frames and the unloaded RTT of the SIRD-Caladan stack is approximately 18 $\upmu{}$s. Latency measurements are end-to-end, measured by the client, and include four request/reply copies to/from the application layer. We sample other system metrics every 2ms. While the size of requests may vary from experiment to experiment, replies are of minimal size.
We configure SIRD parameters as follows: $BDP=216KB$ (24 jumbo frames), $B=1.5\times BDP$, $SThr=0.5\times BDP$, $UnschT=1.0\times BDP$. We do not use switch priority queues.
6.1.1 Receiver Congestion
SIRD’s main objectives when handling incast are eliminating queuing-induced latency for small messages and being able to reduce latency for larger messages through credit-based scheduling. To evaluate this, we run an experiment where six senders saturate a receiver by sending 10MB requests at a rate of 17Gbps each. A seventh sender periodically transmits either 8B or 500KB requests and captures the latency distribution.
Figure 3: Incast: CDF of message latency under incast compared to an unloaded baseline for 8B requests (left) and 500KB requests (right). The incast is formed by six senders transmitting 10MB messages in an open loop.
Figure 3 (left) plots the round-trip latency experienced by short requests, which are unscheduled. When the receiver is saturated, we observe only a few microseconds of additional latency compared to when the receiver is unloaded, which corresponds to a couple of packets of queuing, either at the ToR or at the host stack. For context, the median latency under load for the same experiment using kernel TCP Cubic was above 1ms. SIRD achieves this result because it limits the number of scheduled inbound bytes to $B$ and thus the extent of downlink queuing to $B-BDP$. Further, since in this case senders respond to credit immediately, credit pacing effectively eliminates downlink queuing, while still achieving 96Gbps in this experiment.
Figure 3 (right) shows the latency of scheduled 500KB requests under two receiver scheduling policies. Under SRPT, the receiver allocates credit to the message with the fewest remaining bytes and thus quickly prioritizes 500KB over 10MB requests, leading to near-unloaded latency despite the saturated downlink. SIRD can also implement fairer policies like per-sender round robin ("SRR") as shown. Finally, as it causes minimal network queuing, SIRD can implement such policies without in-network QoS support.
6.1.2 Sender Information
When senders are the bottleneck (outcast), SIRD tackles the congested sender problem by providing congestion feedback to receivers and allowing them to adjust their credit allocation. To evaluate the mechanism’s efficacy at 100Gbps, we run an outcast experiment where a single sender sends 10MB messages at full rate to three receivers in a time-staggered manner. Figure 4 (left) shows that, without informed overcommitment ("$SThr=Inf$"), the level of credit accumulated at the congested sender increases with every new receiver. This is because each receiver independently gives the sender $1\times BDP$ worth of credit to achieve full rate. Through congestion feedback ("$SThr=0.5\times BDP$"), receivers coordinate and scale down their credit allocation until credit accumulation at the sender is less than $SThr=0.5\times BDP$, on average. Credit, instead, stays at receivers, and can be used to schedule other senders. Figure 4 (right) shows the sum of credit available at the three receivers, each receiving a max-min fair share of the sender’s bandwidth. Without informed overcommitment, each receiver allocates $1\times BDP$ worth of credit to the congested sender. With it, each allocates $(BDP+SThr)/num\_rcvers$.
Figure 4: Left: credit accumulated at congested sender. Right: sum of credit available at the three receivers - initial total: $4\times 1.5=4.5\times BDP$. 100ms moving average. The circled numbers indicate the number of receivers at that stage of the experiment.
6.2 Simulations
We use network simulation to evaluate SIRD on large-scale workloads and compare it to other protocol designs.
@@ 6.2 Simulations
We compare SIRD to 5 baselines: DCTCP [8], a widely deployed sender-driven scheme [33], Swift [44], a state-of-the-art production sender-driven scheme, Homa [56], because of its near-optimal latency and its use of overcommitment, ExpressPass [20], as it employs a hop-by-hop approach to managing credit, and dcPIM [16], as a unique point in the design space as it explicitly matches senders and receivers. We do not extend Homa with Aeolus [39] for the reasons discussed in [40]. Note that the Homa and dcPIM papers [56, 16] already include favorable comparisons to NDP [36], Aeolus [39], PIAS [13], pHost [31], and HPCC [47], thus we did not include them in our evaluation.
@@
We implement SIRD on ns-2 [11], reuse the original ns-2 DCTCP implementation (using the same parameters as [10], scaled to 100Gbps) and the original ns-2 ExpressPass implementation [21], and port the published Homa simulator [55] to ns-2 using the same parameters as in [56] scaled to 100Gbps. The published Homa simulator does not implement the incast optimization [56], which further relies on two-way messages. We use a community Swift simulator, kindly made public [3], and configure its delay parameters to achieve similar throughput to DCTCP, respecting the guidelines [44]. Finally, we use the published dcPIM simulator [15]. SIRD’s simulator is open source and available on GitHub [2].
Table 2 lists protocol parameter values. DCTCP and Swift use pools of pre-established connections (40 for each host pair). SIRD approximates SRPT scheduling like Homa and dcPIM. SIRD, Homa, and dcPIM use packet spraying, DCTCP and Swift use ECMP. Homa uses 8 network priority levels, dcPIM 3, and SIRD 2 (§4.4). The initial window of DCTCP and Swift is configured at $BDP$ like in [10]. The RTT latency is similar to prior work.
Topology: We simulate the two-tier leaf-spine topology used in previous work [56, 8, 10, 31, 16] with 144 hosts, connected to 9 top-of-rack switches (16 hosts each) via 4 spine switches, with link speeds of 100Gbps to hosts and 400Gbps to spines. We simulate switches with infinite buffers, i.e., without packet drops (1) to avoid making methodologically complex assumptions as drop rates and thus latency and throughput are very sensitive to switch buffer sizes, organization, and configuration [33, 40] (see  §2.2); and (2) to study the intended mode of operation of the protocols that leverage buffering to achieve high link utilization and operate best without drops. The intentional dropping of credit packets to adjust sender rates by ExpressPass is maintained. SIRD never uses more than a small fraction of the theoretical capacity (§6.2.1) and is not affected by this setup.
Workload: Each host operates both as client and server, sending one-way messages according to an open-loop Poisson distribution to uniformly random receivers (all-to-all). We simulate 3 workloads: (1) WKa: an aggregate of RPC sizes at a Google datacenter [28]; (2) WKb: a Hadoop workload at Facebook [64]; (3) WKc: a web search application [10]. We select them to test over a wide range of mean message sizes of 3KB, 125KB, and 2.5MB respectively.
Table 2: Default simulation parameters for each protocol.
@@
Homa Same as [56] at 100Gbps (incl. priority split)
dcPIM Same as [16]
We simulate these 3 workloads on 3 traffic configurations, for a total of 9 points of comparison. The configurations are: (1) Balanced: The default configuration described above. We vary the applied load, which does not include protocol-dependent header overheads, from 25% to 95% of link capacity. (2) Core: Same as (1) but ToR-Spine links are 200Gbps (2-to-1 oversubscription). Due to uniform message target selection, $128/144\approx 89\%$ of messages travel via spines, turning the core into the bottleneck. We consequently reduce the load applied by hosts by ×$\times 0.89*2$ to reflect the network’s reduced capacity. This is not meant to reflect a permanent load distribution, but we hypothesize that it is possible transiently. (3) Incast: We use the methodology of [47, 16] and combine background traffic with overlay incast traffic: 30 random senders periodically send a 500KB message to a random receiver. Incast traffic represents 7% of the total load.
We report goodput (rate of received application payload), total buffering in switches, and message slowdown, defined as the ratio between the measured and the minimum possible latency for each message. In the incast configuration, we exclude incast messages from slowdown results.
6.2.1 Performance Overview
@@
Figure 5: Normalized goodput, queuing, and slowdown across all 9 configurations. Each metric is normalized based on the best-performing protocol for the given metric and configuration. For queuing and slowdown, lower is better. For goodput, higher is better. Normalized slowdown and buffering are capped at 10×10$\times$10 × and 200×200$\times$200 ×, respectively, and higher values are plotted in the overflow area. The numbers in parentheses show the number of unstable configurations for each protocol which are not plotted. X-axis jitter is added for visibility. Find the data in Appendix A.
@@
Figure 5 shows how protocols navigate the trade-offs between throughput, buffering, and latency by plotting their relative performance across all 9 workload-configuration combinations. The best-performing protocol on each of the 9 scenarios gets a score of $1.0$ (per metric) and the others are normalized to it, so that goodput is always $\leq 1.0$ whereas queuing and slowdown are always $\geq 1.0$. We report the highest achieved goodput and peak queuing over all the levels of applied load. We report $99^{th}$ percentile slowdown across all messages of each workload at 50% applied load which is a level most protocols can deliver in all scenarios. The following figures and calculations do not include cases where a protocol cannot satisfy a specific load level or cannot stop network buffers from growing infinitely.
@@
Overall, SIRD is the only protocol that consistently achieves near-ideal scores across all metrics. Specifically, SIRD causes 12×12$\times$12 × less peak network queuing than Homa and achieves competitive latency and goodput performance. SIRD outperforms dcPIM in message slowdown, and peak goodput and queuing by $46\%$, $9\%$, and $43\%$ respectively. ExpressPass causes practically zero queuing thanks to its hop-by-hop design, and 88% less than SIRD, but SIRD delivers 10×10$\times$10 × lower slowdown and $26\%$ more goodput. Even under full fabric saturation, SIRD induces at most 0.8MB of ToR queuing in receiver-bottleneck scenarios and 2.3MB in core-bottleneck scenarios. Given a packet buffer capacity of 3.13MB/Tbps (see §2.2), these values correspond to a maximum buffer occupancy of 8% and 23%, respectively.
6.2.2 Congestion Response